OpenStudy (anonymous):

write the standard form of the equation of each line given the slope and y-intercept. slop=-3/5, y-intercept=5

OpenStudy (anonymous):

-3/5 times X + 5

OpenStudy (anonymous):

y=-3/5 x+5

OpenStudy (anonymous):

so its result is 3x+5y=5?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

3x+5y=25

OpenStudy (anonymous):

don't know what u mean... standard form is y=mx+b.

OpenStudy (anonymous):

because your result is different with them

OpenStudy (anonymous):

their result is 3x+5y+25

OpenStudy (anonymous):

3x+5y=25